Grade 11Mechanics

A particle starts from rest at A and moves with uniform acceleration a metre per second square in a straight line. after (1/ a) seconds a second particle starts from A and moves with uniform velocity u in the same line and same direction. if u is greater than 2 metre per second then during the entire motion second particle remains ahead of the first particle for a duration?

Ayush
8 Years agoGrade 11

1 Answer

Khimraj
8 Years ago
distance covered by first particle
d1 = ½*a*t2
and distance covered by second particle
d2 = 2(t – 1/a)
and d1 = d2
½*a*t2 = 2(t – 1/a)
at2 = 4t -4/a
a2t2 – 4at + 4=0
(at -2)2 = 0
then t = 2/a